# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ Appendix G Generalize to Stochastic Rewards
$\bar{\mathcal{G}}=\bigcup_{(s_{g},a_{g})\in\mathcal{S}_{g}\times\mathcal{A}_{g}}\mathrm{supp}\left(\mathcal{G}_{s_{g},a_{g}}\right),\quad\quad\bar{\mathcal{L}}=\bigcup_{(s_{i},s_{g},a_{i})\in\mathcal{S}_{l}\times\mathcal{S}_{g}\times\mathcal{A}_{l}}\mathrm{supp}\left(\mathcal{L}_{s_{i},s_{g},a_{i}}\right),$ (40)
@@
Definition G.2.
Let the randomized empirical adapted Bellman operator be $\widehat{\mathcal{T}}^{\text{random}}_{k,m}$ such that:
$\widehat{\mathcal{T}}^{\text{random}}_{k,m}\widehat{Q}^{t}_{k,m}(s_{g},s_{1},F_{z_{\Delta}},a_{g},a_{1})=R_{\Delta}(s,a)+\frac{\gamma}{m}\sum_{\ell\in[m]}\max_{a^{\prime}\in\mathcal{A}}\widehat{Q}_{k,m}^{t}(s_{g}^{\ell},s_{j}^{\ell},F_{z_{\Delta\setminus j}^{\ell}},a_{j}^{\ell},a_{g}^{\ell}),$ (41)
@@
The proposed algorithm averages $\Xi$ samples of the adapted randomized empirical adapted Bellman operator, $\mathcal{T}_{k,m}^{\text{random}}$ and updates the $\widehat{Q}_{k,m}$ function using with the average. One can show that $\mathcal{T}_{k,m}^{\text{random}}$ is a contraction operator with module $\gamma$. By Banach’s fixed point theorem, $\mathcal{T}_{k,m}^{\text{random}}$ admits a unique fixed point $\widehat{Q}_{k,m}^{\text{random}}$.
@@
1:  Set $\widetilde{\Delta}=\{2,\dots,k\}$.
@@
3:  Set $\widehat{Q}^{0}_{k,m}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{g},a_{1})=0$, for $(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{g}\times\mathcal{A}_{l}$.
@@
5:     for $(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{g},a_{1})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}{(\mathcal{Z}_{l})}\times\mathcal{A}_{g}\times\mathcal{A}_{l}$ do
@@
8:           $\rho=\rho+\widehat{\mathcal{T}}^{\text{random}}_{k,m}\widehat{Q}^{t}_{k,m}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{g},a_{1})$
9:        $\widehat{Q}^{t+1}_{k,m}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{g},a_{1})=\rho/\Xi$
@@
$\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta\setminus i}})=\arg\max_{(a_{g},a_{i},F_{a_{\Delta\setminus i}})\in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mu_{k-1}{(\mathcal{Z}_{l})}}\widehat{Q}_{k,m}^{T}(s_{g},s_{i},F_{z_{\Delta\setminus i}},a_{g},a_{i}).$
Theorem G.3 (Hoeffding’s Theorem (Tsybakov, 2008)).
@@
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{\widetilde{\pi}_{k,m}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}\right)\right]\geq 1-\frac{1}{100\sqrt{k}}.$ (43)
@@
Then, applying the triangle inequality to $\epsilon_{k,m}$ allows us to derive a probabilistic bound on the optimality gap between $\widehat{Q}_{k,m}^{\mathrm{est}}$ and $Q^{*}$, where the gap is increased by $\frac{1}{\sqrt{200k}}$, and where the randomness is over the stochasticity of the rewards. Then, the optimality gap between $V^{\pi*}$ and $V^{{\pi}_{k,m}^{\mathrm{est}}}$, for when the policy $\widehat{\pi}^{\mathrm{est}}_{k,m}$ is learned using Algorithm 5 in the presence of stochastic rewards obeying G.1 follows
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{\widetilde{\pi}_{k,m}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}\right)\right]\geq 1-\frac{1}{100\sqrt{k}},$ (46)
@@
and using methods from order statistics to bound the errors of the estimates and use the deviations between estimates to determine an optimal online stopping time (Kleinberg, 2005) as is done in the online secretary problem. This, more sophisticated, argument would become an essential step in converting this to a truly online learning, with a stochastic approximation scheme. Furthermore, one could incorporate variance-based analysis in the algorithm, where we use information derived from higher-order moments to do a weighted update of the Bellman update (Jin et al., 2024), thereby taking advantage of the skewedness of the distribution, and allowing us the freedom to assign an optimism/pessimism score to our estimate of the reward.
Appendix H Extension to Continuous State/Action Spaces
@@
Theorem H.1 (Baire Category Theorem (Jin et al., 2020)).
@@
$\mathbb{P}(\cdot|s,a)=\langle\phi(s,a),\mu(\cdot)\rangle,\quad\quad r(s,a)=\langle\phi(s,a),\theta\rangle$ (47)
Without loss of generality, assume $\|\phi(s,a)\|\leq 1,\forall(s,a)\in S\times A$ and $\max\{\|\mu(S)\|,\|\theta\|\}\leq\sqrt{d}$.
@@
Assumption H.3.
Suppose $\mathcal{S}_{g}\subset\mathbb{R}^{\sigma_{g}},\mathcal{S}_{l}\subset\mathbb{R}^{\sigma_{l}},\mathcal{A}_{g}\subset\mathbb{R}^{\alpha_{g}},\mathcal{A}_{l}\subset\mathbb{R}^{\alpha_{l}}$ are bounded compact sets. From the Baire category theorem, the underlying field $\mathbb{R}$ can be replaced to any set of Baire first category, which satisfies the property that there exists a dense open subset. In particular, replace 2.3 with a boundedness assumption.
These multi-agent settings where each agent handles a continuous state/action space find many applications in optimization, control, and synchronization.
Example H.4 (Federated Learning).
Consider a peer-to-peer learning setting (Chaudhari et al., 2024) with $n$ agents, where each agent possesses a common neural network architecture $f(\mathbf{x};\mathbf{\theta}):\mathbb{R}^{N}\to\mathbb{R}^{M}$, parameterized by $\theta\in\mathbb{R}^{P}$. Here, each local agent has a common local objective function $L_{q}(\mathbf{\theta}):\mathbb{R}^{P}\to\mathbb{R}$ defined using a local dataset $\mathcal{D}_{q}\stackrel{{\scriptstyle\Delta}}{{=}}\{x_{q_{i}},y_{q,i}\}_{i=1}^{D}$. For instance, $L_{q}(\mathbf{\theta})$ can be the local empirical risk function that evaluates the performance of $f(\mathbf{x};\mathbf{\theta})$ on a task involving the local dataset $\mathcal{D}_{q}$, where the agents cooperate to learn the optimal $\mathbf{\theta}^{*}$ that minimizes the global objective, which is a weighted average of all local objectives. Formally, the agents collaboratively aim to find $\mathbf{\theta}^{*}$ satisfying:
$\theta^{*}=\arg\min_{\theta\in\mathbb{R}^{P}}L(\theta)$
$L(\theta)=\frac{1}{Q}\sum_{q=1}^{Q}L_{q}(\theta)\stackrel{{\scriptstyle\Delta}}{{=}}\frac{1}{Q}\sum_{q=1}^{Q}\frac{1}{D}\sum_{i=1}^{D}\ell(f(x_{q,i};\theta),y_{q,i})$
Here, the global agent may act to balance the loss among agents, by ensuring that the variance on the losses across the agents is small. If the variance on the losses of the agents is large, it could assign a large cost to the system, and signal that the policies of the local agents must favor the mean state, promoting convergence.
Example H.5 (Quadcopter Swarm Drone (Preiss et al., 2017)).
Lemma H.6 (Proposition 2.3 in Jin et al. (2020)).
@@
$\phi(s,a)^{\top}\mu(\mathcal{S})=1,\quad\quad\phi(s,a)^{\top}\mu(\mathcal{B})\geq 0.$
Property H.8.
@@
$P_{g}(s_{g}^{\prime}|s_{g},a_{g})=\phi_{g}(s_{g},a_{g})^{\top}\mu_{g}(s_{g}^{\prime})$ (48)
$P_{l}(s_{i}^{\prime}|s_{i},s_{g},a_{i})=\phi_{l}(s_{i},s_{g},a_{i})^{\top}\mu_{l}(s_{i}^{\prime})$ (49)
for some features $\phi_{g}(s_{g},a_{g})\in\mathbb{R}^{d}$ and $\phi_{l}(s_{i},s_{g},a_{i})\in\mathbb{R}^{d}$. Then, the dynamics are amenable to the following spectral factorization, as in Ren et al. (2024).
Lemma H.9.
$\widehat{Q}_{k}$ admits a linear representation.
@@
$=\phi_{g}(s_{g},a_{g}),\mu_{g}(s_{g}^{\prime})\rangle\cdot\left\langle\otimes_{i=1}^{n}\phi_{l}(s_{i},s_{g},a_{i}),\otimes_{i=1}^{n}\mu_{l}(s_{i}^{\prime})\right\rangle\coloneq\langle\bar{\phi}_{n}(s,a),\bar{\mu}_{n}(s^{\prime})\rangle$
@@
Therefore, $\widehat{Q}_{k}$ admits the linear representation:
$\widehat{Q}_{k}^{\pi_{k}}(s_{g},F_{z_{\Delta}},a_{g})$ $=r_{g}(s_{g},a_{g})+\frac{1}{k}\sum_{i\in\Delta}r_{l}(s_{i},s_{g},a_{i})+\gamma\mathbb{E}_{s_{g}^{\prime},s_{\Delta^{\prime}}}\max_{a_{g}^{\prime},a_{\Delta}^{\prime}}\widehat{Q}_{k}^{\pi_{k}}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})$
@@
Therefore, $\bar{\phi}_{k}$ serves as a good representation for the $\widehat{Q}_{k}$-function making the problem amenable to the classic linear functional approximation algorithms, consisting of feature generation and policy gradients.
In feature generation, we generate the appropriate features $\phi$, comprising of the local reward functions and the spectral features coming from the factorization of the dynamics. In applying the policy gradient, we perform a gradient step to update the local policy weights $\theta_{i}$ and update the new policy. For this, we update the weight $w$ via the TD(0) target semi-gradient method (with step size $\alpha$), to get

### v3

@@ Appendix H Generalization to Stochastic Rewards
$\bar{\mathcal{G}}$ $=\bigcup_{(s_{g},a_{g})\in\mathcal{S}_{g}\times\mathcal{A}_{g}}\mathrm{supp}\left(\mathcal{G}_{s_{g},a_{g}}\right)$
$\bar{\mathcal{L}}$ $=\bigcup_{(s_{i},s_{g},a_{i})\in\mathcal{S}_{l}\times\mathcal{S}_{g}\times\mathcal{A}_{l}}\mathrm{supp}\left(\mathcal{L}_{s_{i},s_{g},a_{i}}\right)$
@@
Definition H.2.
Let the randomized empirical adapted Bellman operator be $\hat{\mathcal{T}}^{\text{random}}_{k,m}$ such that:
$\hat{\mathcal{T}}^{\text{random}}_{k,m}\hat{Q}^{t}_{k,m}(s_{g},s_{1},F_{z_{\Delta}},a_{g},a_{1})=R_{\Delta}(s,a)+\frac{\gamma}{m}\sum_{\ell\in[m]}\max_{a^{\prime}\in\mathcal{A}}\hat{Q}_{k,m}^{t}(s_{g}^{\ell},s_{j}^{\ell},F_{z_{\Delta\setminus j}^{\ell}},a_{j}^{\ell},a_{g}^{\ell}),$ (37)
@@
The proposed algorithm averages $\Xi$ samples of the adapted randomized empirical adapted Bellman operator, $\mathcal{T}_{k,m}^{\text{random}}$ and updates the $\hat{Q}_{k,m}$ function using with the average. One can show that $\mathcal{T}_{k,m}^{\text{random}}$ is a contraction operator with module $\gamma$. By Banach’s fixed point theorem, $\mathcal{T}_{k,m}^{\text{random}}$ admits a unique fixed point $\hat{Q}_{k,m}^{\text{random}}$.
@@
1:  Set $\tilde{\Delta}=\{2,\dots,k\}$.
@@
3:  Set $\hat{Q}^{0}_{k,m}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{g},a_{1})=0$, for $(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{g}\times\mathcal{A}_{l}$.
@@
5:     for $(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{g},a_{1})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}{(\mathcal{Z}_{l})}\times\mathcal{A}_{g}\times\mathcal{A}_{l}$ do
@@
8:           $\rho=\rho+\hat{\mathcal{T}}^{\text{random}}_{k,m}\hat{Q}^{t}_{k,m}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{g},a_{1})$
9:        $\hat{Q}^{t+1}_{k,m}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{g},a_{1})=\rho/\Xi$
@@
$\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta\setminus i}})=\mathop{\operatorname*{arg\,max}}_{(a_{g},a_{i},F_{a_{\Delta\setminus i}})\in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mu_{k-1}{(\mathcal{Z}_{l})}}\hat{Q}_{k,m}^{T}(s_{g},s_{i},F_{z_{\Delta\setminus i}},a_{g},a_{i}).$
Theorem H.3 (Hoeffding’s Theorem [Tsybakov, 2008]).
@@
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{\tilde{\pi}_{k,m}}(s_{0})\leq\tilde{O}\left(\frac{1}{\sqrt{k}}\right)\right]\geq 1-\frac{1}{100\sqrt{k}}.$ (39)
@@
Then, applying the triangle inequality to $\epsilon_{k,m}$ allows us to derive a probabilistic bound on the optimality gap between $\hat{Q}_{k,m}^{\mathrm{est}}$ and $Q^{*}$, where the gap is increased by $\frac{1}{\sqrt{200k}}$, and where the randomness is over the stochasticity of the rewards. Then, the optimality gap between $V^{\pi*}$ and $V^{{\pi}_{k,m}^{\mathrm{est}}}$, for when the policy $\hat{\pi}^{\mathrm{est}}_{k,m}$ is learned using Algorithm 4 in the presence of stochastic rewards obeying H.1 follows
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{\tilde{\pi}_{k,m}}(s_{0})\leq\tilde{O}\left(\frac{1}{\sqrt{k}}\right)\right]\geq 1-\frac{1}{100\sqrt{k}},$ (41)
@@
and using methods from order statistics to bound the errors of the estimates and use the deviations between estimates to determine an optimal online stopping time [Kleinberg, 2005] as is done in the online secretary problem. This, more sophisticated, argument would become an essential step in converting this to a truly online learning, with a stochastic approximation scheme. Furthermore, one could incorporate variance-based analysis in the algorithm, where we use information derived from higher-order moments to do a weighted update of the Bellman update [Jin et al., 2024], thereby taking advantage of the skewedness of the distribution, and allowing us the freedom to assign an optimism/pessimism score to our estimate of the reward.
Appendix I Partially Relaxing the Offline Learning assumption: Off-Policy Learning
A limitation of the planning Algorithm 1 is that it learns $\hat{Q}_{k}^{*}$ in an offline manner by assuming a generative oracle access to the transition functions $\mathbb{P}_{g},\mathbb{P}_{l}$, and reward function $r(\cdot,\cdot)$. In certain realistic RL applications, such a generative oracle might not exist, and it is more desirable to perform off-policy learning where the agent continues to learn in an offline manner but from historical data Fujimoto et al. [2019]. In this setting, the agents learn the target policy $\hat{\pi}_{k}^{*}$ using data generated by a different behavior policy (the strategy it uses to explore the environment). There is a significant body of work on the theoretical guarantees in off-policy learning Chen et al. [2021b], Chen and Maguluri [2022b], Chen et al. [2021a, 2025]. In fact, these previous results are amenable to transforming guarantees about offline $Q$-learning to off-policy $Q$-learning, at the cost of $\log|\mathcal{S}_{g}||\mathcal{S}_{l}|^{k}|\mathcal{A}_{g}||\mathcal{A}_{l}|^{k}$ factors in the runtime. Therefore, this section is devoted to showing that our previous result satisfy the assumptions of transforming offline $Q$-learning to off-policy $Q$-learning for the subsampled $\hat{Q}_{k}$-function, and we further show that, in expectation, can maintain the decaying optimality gap of $\tilde{O}(1/\sqrt{k})$ of the learned policy $\pi_{k}$, where the randomness is over the heuristic exploration policy $\pi_{b}$.
The off-policy $\hat{Q}_{k}$-learning algorithm is an iterative algorithm to estimate the optimal $\hat{Q}_{k}$-function as follows: first, a sample trajectory $\{(s_{g},s_{\Delta},a_{g},a_{\Delta})\}$ is collected using a suitable behavior policy $\pi_{k,b}$. For simplicity of notation, let $S_{\Delta}=(s_{g},s_{\Delta})$ and $A_{\Delta}=(a_{g},a_{\Delta})$. Then, initialize $\hat{Q}_{k}^{0}:{|\mathcal{S}_{g}||\mathcal{S}_{l}|^{k}|\mathcal{A}_{g}||\mathcal{A}_{l}|^{k}}\to\mathbb{R}$ and let $\alpha>0$ be determined later. For each $t\geq 0$ and state-action pair $(S_{\Delta},A_{\Delta})$ that is updated to $S_{\Delta}^{\prime}$, the iterate $\hat{Q}_{k}^{t}(S_{\Delta},A_{\Delta})$ is updated by
$\hat{Q}_{k}^{t+1}(S_{\Delta},A_{\Delta})=(1-\alpha)\hat{Q}_{k}^{t}(S_{\Delta},A_{\Delta})+\alpha_{t}\left(r(S_{\Delta},A_{\Delta})+\gamma\max_{A_{\Delta}^{\prime}\in\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}}\hat{Q}_{k}^{t}(S_{\Delta}^{\prime},A_{\Delta}^{\prime})\right).$ (42)
Note that the update in Equation 42 does not include an expectation and can be computed in a single trajectory via historical data. We make the following assumption:
Assumption I.1.
The behavior policy $\pi_{b}$ satisfies $\pi_{b}(A_{\Delta}|S_{\Delta})>0$ for all $(S_{\Delta},A_{\Delta})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$ and the Markov chain $\mathcal{M}_{S_{\Delta}}=\{S_{\Delta}^{(t)}\}_{t\geq 0}$ induced by $\pi_{b}$ is irreducible and aperiodic with stationary distribution $\mu$ and mixing time $t_{\delta}(\mathcal{M})=\min\{t\geq 0:\max_{S_{\Delta}\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}}\|P^{t}(S_{\Delta},\cdot)-\mu(\cdot)\|_{TV}\leq\delta\}$. There are many heuristics of such behavior policies [Fujimoto et al., 2019].
Theorem I.2.
Let $\pi_{k}$ be the policy learned through off-policy $\hat{Q}_{k}$-learning. Then, under I.1, we have that that with probability at least $1-\frac{1}{100e^{k}}$,
$\mathbb{E}[V^{\pi^{*}}(s_{0})-V^{\pi_{k}}(s_{0})]$ $\leq\frac{\tilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{40\tilde{r}|\mathcal{S}_{l}||\mathcal{A}_{l}||\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|+\frac{1}{2}}}{(1-\gamma)^{2}}}+\frac{1}{10\sqrt{k}}+2\epsilon_{k,m}$
$=\tilde{O}(1/\sqrt{k}),$
where the randomness in the expectation is over the stochasticity of the exploration policy $\pi_{b}$.
Proposition I.3.
Recall that the following are true:
  1. 1.
    $\|\hat{Q}_{k}(F_{S_{\Delta}},F_{A_{\Delta}})-\hat{Q}_{k}(F_{S^{\prime}_{\Delta}},F_{A_{\Delta}})\|\leq\frac{1}{1-\gamma}\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\|F_{S^{\prime}_{\Delta}}-F_{S_{\Delta}}\|_{1}$, for any $S_{\Delta},S^{\prime}_{\Delta}\in\mathcal{S}_{g}\times\mathcal{A}_{l}^{k}$ and $A_{\Delta}\in\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$ by Theorem E.3,
  2. 2.
    $\|\hat{Q}_{k}\|\leq\frac{\tilde{r}}{1-\gamma}$ by Lemma C.4,
  3. 3.
    $\|\hat{Q}^{t+1}_{k}(S_{\Delta},A_{\Delta})-\tilde{Q}_{k}^{t+1}(S_{\Delta},A_{\Delta})\|_{\infty}\leq\gamma\|\hat{Q}_{k}^{t}\ -\tilde{Q}_{k}^{t}(S_{\Delta},A_{\Delta})\|_{\infty}$ by Lemma C.6,
  4. 4.
    The Markov chain $\mathcal{M}_{S}$ enjoys a rapid mixing property from I.1,
Then, by treating the single trajectory update of the $\hat{Q}_{k}$-function as a noisy addition to the expected update from the ideal Bellman operator, Chen et al. [2021b] uses Markovian stochastic approximation to bound $\mathbb{E}[\|\hat{Q}_{k}^{T}-\hat{Q}_{k}^{*}\|_{\infty}^{2}]$. We restate their result:
Theorem I.4 (Theorem 3.1 in Chen et al. [2021b] adapted to our setting).
Suppose $\alpha_{t}=\alpha$ for all $t\geq 0$, where $\alpha$ is chosen such that $\alpha t_{\alpha}(\mathcal{M}_{S_{\Delta}})\leq c_{Q,0}\frac{(1-\beta_{1})^{2}}{\log|\mathcal{S}_{g}||\mathcal{S}_{l}^{k}||\mathcal{A}_{g}|\mathcal{A}_{l}^{k}|}$, where $c_{Q,0}$ is a numerical constant. Then, under Proposition I.3, for all $t\geq t_{\alpha}(\mathcal{M}_{S_{\Delta}})$, we have
$\mathbb{E}[\|\hat{Q}^{t}_{k}-\hat{Q}_{k}^{*}\|^{2}_{\infty}]\leq c_{Q,1}\left(1-\frac{(1-\gamma)\alpha}{2}\right)^{t-t_{\alpha}(\mathcal{M}_{S_{\Delta}})}+c_{Q,2}\frac{\log|\mathcal{S}_{g}||\mathcal{S}_{l}|^{k}|\mathcal{A}_{g}||\mathcal{A}_{l}|^{k}}{(1-\gamma)^{2}}\alpha t_{\alpha}(\mathcal{M}_{S_{\Delta}}),$
where $c_{Q,1}=3(\frac{\tilde{r}}{1-\gamma}+1)^{2}$, $c_{Q,2}=912e(\frac{3\tilde{r}}{1-\gamma}+1)^{2}$, and where the randomness in the expectation is over the randomness of the stochasticity of the behavior/exploration policy $\pi_{b}$.
Corollary I.5 (Corollary 3.2 in Chen et al. [2021b] adapted to our setting).
To make $\mathbb{E}[\|\hat{Q}_{k}^{t}-\hat{Q}_{k}^{*}\|_{\infty}\leq\frac{1}{100\sqrt{k}}]$ for $\epsilon>0$, we need
$t>\tilde{O}(\frac{10000k\log^{2}(100\sqrt{k})|\mathcal{S}_{g}||\mathcal{S}_{l}|^{k}|\mathcal{A}_{g}||\mathcal{A}_{l}|^{k}\log|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{S}_{l}|^{k}|\mathcal{A}_{l}|^{k}}{(1-\gamma)^{5}}).$
With this sample complexity, by the triangle inequality we also recover an expected-value analog of Theorem F.8.
Corollary I.6.
For $\delta\in(0,1)^{2}$, with probability at least $1-\delta$, we have
$\mathbb{E}[\hat{Q}_{k}^{*}(s_{g},F_{s_{\Delta}},a_{g},F_{a_{\Delta}})-Q_{n}^{*}(s_{g},F_{s_{[n]},a_{g},F_{a_{[n]}}})]\leq\frac{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}{1-\gamma}\sqrt{\frac{n-k+1}{8kn}}\|r_{l}\|_{\infty},$
where the randomness in the expectation is over the stochasticity of the exploration policy $\pi_{b}$.
In turn, following the argument in the proof of Theorem G.3, it is straightforward to verify that this leads to a result on the expected performance difference using off-policy learning:
Corollary I.7.
With probability at least $1-\frac{1}{100e^{k}}$,
$\mathbb{E}[V^{\pi^{*}}(s_{0})-V^{\pi_{k}}(s_{0})]$ $\leq\frac{\tilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{40\tilde{r}|\mathcal{S}_{l}||\mathcal{A}_{l}||\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|+\frac{1}{2}}}{(1-\gamma)^{2}}}+\frac{1}{10\sqrt{k}}+2\epsilon_{k,m}$
$=\tilde{O}(1/\sqrt{k}),$
where the randomness in the expectation is over the stochasticity of the exploration policy $\pi_{b}$.
Appendix J Extension to Continuous State/Action Spaces
Multi-agent settings where each agent handles a continuous state and action space find many applications in optimization, control, and synchronization.
Example J.1 (Quadcopter Swarm Drone [Preiss et al., 2017]).
Hence, this section is devoted to extending our algorithm and theoretical results to the case where the state and action spaces can be continuous (and therefore have infinite cardinality).
@@ Appendix I Partially Relaxing the Offline Learning assumption: Off-Policy Learning
Theorem J.2 (Baire Category Theorem [Jin et al., 2020]).
@@
$\mathbb{P}(\cdot|s,a)$ $=\langle\phi(s,a),\mu(\cdot)\rangle,$
$r(s,a)$ $=\langle\phi(s,a),\theta\rangle$
Assumption J.4 (Bounded features).
Without loss of generality, we assume boundedness of the features: $\|\phi(s,a)\|\leq 1$, for all $(s,a)\in S\times A$ and $\max\{\|\mu(S)\|,\|\theta\|\}\leq\sqrt{d}$.
@@
Then, through a reduction from Ren et al. [2024] that uses function approximation to learn the spectral features $\phi_{k}$ for $\hat{Q}_{k}$, we derive a performance guarantee for the learned policy $\pi_{k}^{\mathrm{est}}$, where the optimality gap decays with $k$:
Theorem J.5.
When ${\pi}^{\mathrm{est}}_{k}$ is derived from the spectral features $\phi_{k}$ learned in $\hat{Q}_{k}$, and $M$ is the number of samples used in the function approximation, then
$\Pr\bigg{[}V^{\pi^{*}}(s)-V^{{\pi}^{\mathrm{est}}_{k}}(s)$ $\leq\tilde{O}\bigg{(}\frac{1}{\sqrt{k}}+\frac{\|{\phi}_{k}\|^{5}\log 2k^{2}}{\sqrt{M}}+\frac{2\gamma\tilde{r}}{(1-\gamma)\sqrt{k}}\|\phi_{k}\|\bigg{)}\bigg{]}\geq 1+\frac{1}{50k}-\frac{201}{100\sqrt{k}}$
We assume the system is a linear MDP, where $\mathcal{S}_{g}$ and $\mathcal{S}_{l}$ are infinite compact sets. By a reduction from Ren et al. [2024] and using function approximation to learn the spectral features $\phi_{k}$ for $\hat{Q}_{k}$, we derive a performance guarantee for the learned policy $\pi_{k}^{\mathrm{est}}$, where the optimality gap decays with $k$.
Assumption J.6.
Suppose $\mathcal{S}_{g}\subset\mathbb{R}^{\sigma_{g}},\mathcal{S}_{l}\subset\mathbb{R}^{\sigma_{l}},\mathcal{A}_{g}\subset\mathbb{R}^{\alpha_{g}},\mathcal{A}_{l}\subset\mathbb{R}^{\alpha_{l}}$ are bounded compact sets. From the Baire category theorem, the underlying field $\mathbb{R}$ can be replaced to any set of Baire first category, which satisfies the property that there exists a dense open subset. In particular, replace 2.1 with a boundedness assumption.
Multi-agent settings where each agent handles a continuous state/action space find many applications in optimization, control, and synchronization.
Lemma J.7 (Proposition 2.3 in Jin et al. [2020]).
@@
$\phi(s,a)^{\top}\mu(\mathcal{S})$ $=1,$
$\phi(s,a)^{\top}\mu(\mathcal{B})$ $\geq 0.$
Property J.9.
@@
$P_{g}(s_{g}^{\prime}|s_{g},a_{g})$ $=\phi_{g}(s_{g},a_{g})^{\top}\mu_{g}(s_{g}^{\prime})$
$P_{l}(s_{i}^{\prime}|s_{i},s_{g},a_{i})$ $=\phi_{l}(s_{i},s_{g},a_{i})^{\top}\mu_{l}(s_{i}^{\prime})$
for some features $\phi_{g}(s_{g},a_{g})\in\mathbb{R}^{d}$ and $\phi_{l}(s_{i},s_{g},a_{i})\in\mathbb{R}^{d}$. Then, the dynamics are amenable to the following spectral factorization, as in Ren et al. [2024].
Lemma J.10.
$\hat{Q}_{k}$ admits a linear representation.
@@
$=\phi_{g}(s_{g},a_{g}),\mu_{g}(s_{g}^{\prime})\rangle\cdot\left\langle\otimes_{i=1}^{n}\phi_{l}(s_{i},s_{g},a_{i}),\otimes_{i=1}^{n}\mu_{l}(s_{i}^{\prime})\right\rangle$
:-$\coloneq\langle\bar{\phi}_{n}(s,a),\bar{\mu}_{n}(s^{\prime})\rangle:$
@@
Therefore, $\hat{Q}_{k}$ admits the linear representation:
$\hat{Q}_{k}^{\pi_{k}}(s_{g},F_{z_{\Delta}},a_{g})$ $=r_{g}(s_{g},a_{g})+\frac{1}{k}\sum_{i\in\Delta}r_{l}(s_{i},s_{g},a_{i})+\gamma\mathbb{E}_{s_{g}^{\prime},s_{\Delta^{\prime}}}\max_{a_{g}^{\prime},a_{\Delta}^{\prime}}\hat{Q}_{k}^{\pi_{k}}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})$
@@
Therefore, $\bar{\phi}_{k}$ serves as a good representation for the $\hat{Q}_{k}$-function making the problem amenable to the classic linear functional approximation algorithms, consisting of feature generation and policy gradients.
In feature generation [Ren et al., 2024], we generate the appropriate features $\phi$, comprising of the local reward functions and the spectral features coming from the factorization of the dynamics. In applying the policy gradient, we perform a gradient step to update the local policy weights $\theta_{i}$ and update the new policy.
For this, we update the weight $w$ via the TD(0) target semi-gradient method (with step size $\alpha$), to get